For what value of x is the tangent of the graph of the function f (x) = x4− ax 3−4 parallel to the x axis?
1
Expert's answer
2019-12-10T11:06:31-0500

From the geometric interpretation of derivative it follows that a tangent line to the graph of the function "f(x)" at the point "(x_0,f(x_0))" is parallel to the "x" axis if, and only if "f'(x_0) = 0."

Then, we have the equation

"f'(x) = (x^4 - ax^3 - 4)' = 4x^3 - 3ax^2 = x^2(4x - 3a) = 0"

It has two solutions: "x_1 = 0,\; x_2 = \frac{3}{4}a."


Answer: Either "x=0" or "x = \frac{3}{4}a."